# Monotonicity of the ultrafilter number function

### v3

@@ Abstract.
We investigate whether the ultrafilter number function $\kappa\mapsto\mathfrak{u}(\kappa)$ on the cardinals is monotone, that is, $\mathfrak{u}(\lambda)\leq\mathfrak{u}(\kappa)$ holds for all cardinals $\lambda<\kappa$ or not. We show that monotonicity can fail, but the failure has a large cardinal strength. On the other hand, we prove that there are many restrictions of the failure of monotonicity. For instance, if $\kappa$ is a singular cardinal with countable cofinality or a strong limit singular cardinal, then $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})$ holds.
@@ 1. Introduction
For a cardinal $\kappa$, let $\mathfrak{u}(\kappa)$ denote the ultrafilter number at $\kappa$, that is, the minimal cardinality of a base of a uniform ultrafilter over $\kappa$. A central research on the ultrafilter number is comparison between $\mathfrak{u}(\kappa)$ and other cardinal invariants at $\kappa$: A typical one is, whether $\mathfrak{u}(\kappa)<2^{\kappa}$ is consistent or not. In this context, the ultrafilter number at $\omega$ is studied extensively, and recently the ultrafilter number at singular cardinals is actively researched (e.g., [3], [6], [7], [8], [9], [11]). In contrast, little is known about the ultrafilter number at successor cardinals. While it is still open if $\mathfrak{u}(\omega_{1})<2^{\omega_{1}}$ is consistent or not, Raghavan and Shelah [17] proved the consistency of $\mathfrak{u}(\omega_{\omega+1})<2^{\omega_{\omega+1}}$ under some large cardinal assumption. However it is unknown if this large cardinal assumption is necessary or not.
This paper also investigates the ultrafilter number, but does not aim to study comparison between $\mathfrak{u}(\kappa)$ and other cardinal invariants at $\kappa$. Instead, we identify the ultrafilter numbers as the function $\kappa\mapsto\mathfrak{u}(\kappa)$ on the cardinals, and we shall consider a behavior of the ultrafilter number function. Under GCH, the ultrafilter number $\mathfrak{u}(\kappa)$ is just $2^{\kappa}=\kappa^{+}$, hence the ultrafilter number function is just the continuum function $\kappa\mapsto 2^{\kappa}$, in particular it is strictly increasing. On the other hand, it is easy to construct a model in which the ultrafilter number function is not strictly increasing: Assuming GCH, just add $\omega_{2}$ many Cohen real. Then $\mathfrak{u}(\omega)=\mathfrak{u}(\omega_{1})=\omega_{2}$ holds in the generic extension (e.g., see Lemma 4.1). Under this viewpoint, a natural question arises. This is a main topic of this paper.
@@
For some cardinal invariant, it is easy to show that monotonicity can fail: For the dominating number $\mathfrak{d}(\kappa)$, suppose GCH and add $\omega_{3}$ many Cohen reals. Then $\mathfrak{d}(\omega)=\omega_{3}>\omega_{2}=\mathfrak{d}(\omega_{1})$ holds in the generic extension. However, this naive approach does not work for the ultrafilter numbers; In the generic extension, we have $\mathfrak{u}(\omega)=\mathfrak{u}(\omega_{1})=\omega_{3}$ (see Lemma 4.1 again).
In this paper, we prove that the failure of monotonicity is actually consistent. On the other hand, we also prove that the failure of it has a large cardinal strength. Moreover there are many $\mathsf{ZFC}$-restrictions of the failure of monotonicity. First, by using Raghavan and Shelah’s argument, we show that monotonicity of the ultrafilter number function can fail.
@@
Under certain large cardinal assumptions, the following are consistent:
@@
Raghavan and Shelah’s argument involves indecomposable ultrafilters, but it is not clear if the use of indecomposable ultrafilters is necessary for their theorem. We see that, however, the failure of monotonicity is directly connected with indecomposability. Using this connection, we also prove that there are many restrictions of the failure of monotonicity.
@@
    $\mathfrak{u}(\omega)\leq\mathfrak{u}(\kappa)$ holds for every cardinal $\kappa$.
@@
    For cardinals $\kappa,\lambda$, if $\lambda$ is regular, $\lambda<\kappa$, but $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$, then $\lambda^{+\omega}\leq\kappa$.
@@
    If $\kappa$ is a singular cardinal with countable cofinality, or a strong limit singular cardinal, then $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})$.
@@
    For a singular cardinal $\kappa$, if $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then the set $\{\lambda<\kappa\mid$ $\lambda$ is regular, $\mathfrak{u}(\lambda)>\mathfrak{u}(\kappa^{+})\}$ is bounded in $\kappa$.
@@
    If $\square(\kappa)$ holds for a regular $\kappa$, then $\mathfrak{u}(\lambda)\leq\mathfrak{u}(\kappa)$ holds for every $\lambda<\kappa$.
Finally we show that the failure of monotonicity has a large cardinal strength, so the large cardinal assumption in Theorem 1.2 cannot be eliminated.
Theorem 1.4.
@@
Theorem 1.5.
Suppose there are cardinals $\kappa,\lambda$ such that $\lambda\leq\kappa$ but $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\lambda)$. Then there is an inner models of a proper class of strong cardinals.
@@ 2. Preliminaries
Throughout this paper, a filter means a non-principal proper filter over a non-empty set. A filter $\mathcal{F}$ over a set $S$ is uniform if $\left|{X}\right|=\left|{S}\right|$ for every $X\in\mathcal{F}$. For an ultrafilter $\mathcal{U}$ over a set $S$, a subfamily $\mathcal{G}\subseteq\mathcal{U}$ is a base for $\mathcal{U}$ if for every $Y\in\mathcal{U}$ there is $X\in\mathcal{G}$ with $X\subseteq Y$. Let $\chi(\mathcal{U})=\min\{\left|{\mathcal{G}}\right|\mid\mathcal{G}\subseteq\mathcal{U}$ is a base}}\}}. $\chi(\mathcal{U})$ is called the character of $\mathcal{U}$. Let $\mathfrak{u}(\kappa)$ be the cardinal $\min\{\chi(\mathcal{U})\mid\mathcal{U}$ is a uniform ultrafilter over $\kappa\}$, it is called the ultrafilter number at $\kappa$. It is clear that $\mathfrak{u}(\kappa)\leq 2^{\kappa}$.
Fact 2.1 (Garti and Shelah [6]).
@@
    There is $f:S\to\lambda$ such that for every $X\subseteq\lambda$ with size $<\lambda$, we have $f^{-1}(X)\notin\mathcal{F}$.
@@
Fact 2.4 (Prikry and Silver [16]).
Let $\lambda<\kappa$ be regular cardinals. If $\kappa$ has a $\lambda$-indecomposable uniform ultrafilter, then every stationary subset of $\{\alpha<\kappa\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$ is reflecting, that is, for every stationary $S\subseteq\{\alpha<\kappa\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$, there is $\beta<\kappa$ such that $S\cap\beta$ is stationary in $\beta$.
See also Inamdar and Rinot [12], and Lambie-Hanson and Rinot [15] for recent developments in this direction.
@@
Fact 2.6 (Ben-David and Magidor [2]).
@@
The following lemma might be folklore:
@@
Let $\mathcal{U}$ be an ultrafilter over a set $S$. For a set $T$ and $f:S\to T$, let $f_{*}(\mathcal{U})=\{X\subseteq T\mid f^{-1}(X)\in\mathcal{U}\}$. $f_{*}(\mathcal{U})$ is an ultrafilter (possibly principal) over $T$. It is easy to check that if $\mathcal{U}$ is $\lambda$-indecomposable, then so is $f_{*}(\mathcal{U})$.
The next lemma can be verified easily.
@@
An ultrafilter $\mathcal{U}$ over $S$ is $\lambda$-decomposable if and only if there is a function $f:S\to\lambda$ such that $f_{*}(\mathcal{U})$ is a uniform ultrafilter over $\lambda$.
For a regular cardinal $\kappa$ and a non-empty set $X$, let $\mathrm{Add}(\kappa,X)$ be the poset of all partial functions $p:\kappa\times X\to 2$ with size $<\kappa$. The order is given by the reverse inclusion. The poset $\mathrm{Add}(\kappa,X)$ is $\kappa$-closed, satisfies the $(2^{<\kappa})^{+}$-c.c., and adds $X$-many new subsets of $\kappa$. For $Y\subseteq X$, we will identify $\mathrm{Add}(\kappa,Y)$ as a complete suborder of $\mathrm{Add}(\kappa,X)$.
@@ 3. Raghavan and Shelah’s theorem
Let $\mathcal{N}=\{\dot{A}\mid\dot{A}$ is a $\gamma$-nice name for some $\gamma<\mu\}$. Since $\mathbb{P}$ has the $\mathord{\mathrm{cf}}(\mu)$-c.c., |$\left|{\{p_{i}\mid i<\gamma\}}\right|<\mu$ for $\gamma<\mu$, and $\nu^{\kappa}<\mu$ for every $\nu<\mu$, we have $\left|{\mathcal{N}}\right|\leq\mu$.
@@
We know that the cardinality of $\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega_{1},\mu)$ is just $\mu$, and has the $\omega_{2}$-c.c. By the standard argument, $\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega_{1},\mu)$ forces that $\kappa=2^{\omega}$ and $2^{\kappa}>2^{\omega_{1}}=\mu$. By Proposition 3.1, we have $\mathfrak{u}(\kappa)\leq\mu<2^{\kappa}$ in the generic extension. ∎
The same argument together with Fact 2.6 proves:
@@
In the Prikry forcing extension $V^{\mathbb{P}_{\mathcal{U}}}$, it is known that $\mathcal{U}$ generates an $\nu$-indecomposable uniform ultrafilter for every regular uncountable $\nu<\kappa$ (e.g., see Kanamori [14]). Hence, the forcing with $\mathrm{Add}(\omega,\mu)$ over $V^{\mathbb{P}_{\mathcal{U}}}$ forces that $\kappa$ is singular with countable cofinality and $\mathfrak{u}(\kappa)\leq\mu$ by Proposition 3.1. Moreover $\mu<2^{\kappa}$ since $\mathord{\mathrm{cf}}(\mu)<\kappa$. ∎

### v4

@@ Abstract.
We investigate whether the ultrafilter number function $\kappa\mapsto\mathfrak{u}(\kappa)$ on the cardinals is monotone, that is, whether $\mathfrak{u}(\lambda)\leq\mathfrak{u}(\kappa)$ holds for all cardinals $\lambda<\kappa$ or not. We show that monotonicity can fail, but the failure has large cardinal strength. On the other hand, we prove that there are many restrictions of the failure of monotonicity. For instance, if $\kappa$ is a singular cardinal with countable cofinality or a strong limit singular cardinal, then $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})$ holds.
@@ 1. Introduction
For a cardinal $\kappa$, let $\mathfrak{u}(\kappa)$ denote the ultrafilter number at $\kappa$, that is, the minimal cardinality of a base of a uniform ultrafilter over $\kappa$. A central research on the ultrafilter number is the comparison between $\mathfrak{u}(\kappa)$ and other cardinal invariants at $\kappa$: A typical one is, whether $\mathfrak{u}(\kappa)<2^{\kappa}$ is consistent or not. In this context, the ultrafilter number at $\omega$ is studied extensively, and recently the ultrafilter number at singular cardinals is actively researched (e.g., [4], [8], [9], [10], [11], [13]). In contrast, little is known about the ultrafilter number at successor cardinals. While it is still open if $\mathfrak{u}(\omega_{1})<2^{\omega_{1}}$ is consistent or not, Raghavan and Shelah [21] proved the consistency of $\mathfrak{u}(\omega_{\omega+1})<2^{\omega_{\omega+1}}$ under some large cardinal assumption. However it is unknown if this large cardinal assumption is necessary or not.
This paper also investigates the ultrafilter number, but does not aim to study the comparison between $\mathfrak{u}(\kappa)$ and other cardinal invariants at $\kappa$. Instead, we identify the ultrafilter numbers as the function $\kappa\mapsto\mathfrak{u}(\kappa)$ on the cardinals, and we shall consider the behavior of the ultrafilter number function. Under GCH, the ultrafilter number $\mathfrak{u}(\kappa)$ is just $2^{\kappa}=\kappa^{+}$, hence the ultrafilter number function is just the continuum function $\kappa\mapsto 2^{\kappa}$, in particular it is strictly increasing. On the other hand, it is easy to construct a model in which the ultrafilter number function is not strictly increasing: Assuming GCH, just add $\omega_{2}$ many Cohen real. Then $\mathfrak{u}(\omega)=\mathfrak{u}(\omega_{1})=\omega_{2}$ holds in the generic extension (e.g., see Lemma 4.1). Under this viewpoint, a natural question arises. This is a main topic of this paper.
@@
For some cardinal invariant, it is easy to show that monotonicity can fail: For the dominating number $\mathfrak{d}(\kappa)$, suppose GCH and add $\omega_{3}$ many Cohen reals. Then $\mathfrak{d}(\omega)=\omega_{3}>\omega_{2}=\mathfrak{d}(\omega_{1})$ holds in the generic extension. However, this naive approach does not work for the ultrafilter numbers; In that generic extension, we have $\mathfrak{u}(\omega)=\mathfrak{u}(\omega_{1})=\omega_{3}$ (see Lemma 4.1 again).
In this paper, we prove that the failure of monotonicity is actually consistent. On the other hand, we also prove that the failure of it has large cardinal strength. Moreover there are many $\mathsf{ZFC}$-restrictions of the failure of monotonicity. First, by using Raghavan and Shelah’s argument, we show that monotonicity of the ultrafilter number function can fail.
@@
Relative to a certain large cardinals, the following are consistent [^1] For the exact large cardinal assumption in each item, see the referenced number next to it.:
@@
Raghavan and Shelah’s argument involves indecomposable ultrafilters (see below for indecomposable ultrafilters), but it is not clear if the use of indecomposable ultrafilters is necessary for their theorem. However, as we will see the failure of monotonicity is directly connected with indecomposability. Using this connection, we also prove that there are many restrictions of the failure of monotonicity.
@@
    $\mathfrak{u}(\omega)\leq\mathfrak{u}(\kappa)$ holds for every cardinal $\kappa$ (Proposition 5.3)
@@
    For cardinals $\kappa,\lambda$, if $\lambda$ is regular, $\lambda<\kappa$, but $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$, then $\lambda^{+\omega}\leq\kappa$ (Proposition 5.4).
@@
    If $\kappa$ is a singular cardinal with countable cofinality, or a strong limit singular cardinal, then $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})$ (Corollary 7.5, Theorem 7.9).
@@
    For a singular cardinal $\kappa$, if $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then the set $\{\lambda<\kappa\mid$ $\lambda$ is regular, $\mathfrak{u}(\lambda)>\mathfrak{u}(\kappa^{+})\}$ is bounded in $\kappa$ (Theorem 7.15).
@@
    If $\square(\kappa)$ holds for a regular $\kappa$, then $\mathfrak{u}(\lambda)\leq\mathfrak{u}(\kappa)$ holds for every $\lambda<\kappa$ (Corollary 8.13).
To prove (3) and (4) in this theorem, we will show some new results about (in)decomposable ultrafilters, which are interesting in its own right. Recall that, for a cardinal $\kappa$ and an ultrafilter $\mathcal{U}$ over a set $S$, $\mathcal{U}$ is $\kappa$-decomposable if there is a function $f:S\to\kappa$ such that $f^{-1}(X)\notin\mathcal{U}$ for every $X\in[S]^{<\kappa}$, and it is $\kappa$-indecomposable if it is not $\kappa$-decomposable.
Theorem 1.4 (Theorem 7.3).
Let $\kappa$ be a singular cardinal, and $\mathcal{U}$ an ultrafilter over a set $S$. If $\mathcal{U}$ is $\kappa^{+}$-decomposable and $\mathord{\mathrm{cf}}(\kappa)$-decomposable, then $\mathcal{U}$ is $\kappa$-decomposable as well.
Since every $\kappa$-decomposable ultrafilter is $\mathord{\mathrm{cf}}(\kappa)$-decomposable, this theorem tell us that $\mathord{\mathrm{cf}}(\kappa)$-decomposability and $\kappa$-decomposability are equivalent on $\kappa^{+}$-decomposable ultrafilters.
For a limit cardinal $\kappa$, let us say that an ultrafilter $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable if there is $\lambda<\kappa$ such that $\mathcal{U}$ is $\mu$-decomposable for every regular $\mu$ with $\lambda<\mu<\kappa$. Kunen and Prikry ([17]) showed that, if $\kappa$ is singular, $\mathcal{U}$ is a $\kappa^{+}$-decomposable but $\mathord{\mathrm{cf}}(\kappa)$-indecomposable ultrafilter, then $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable. Lipparini ([19]) showed the converse for $\mathord{\mathrm{cf}}(\kappa)$-indecomposable ultrafilters: If $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable, then $\mathcal{U}$ is $\kappa^{+}$-decomposable if and only if $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable (see Theorem 7.10 and 7.13 below for details). By using Theorem 1.4, we shall prove that $\mathord{\mathrm{cf}}(\kappa)$-indecomposable assumption in their theorem can be replaced by $\kappa$-indecomposability (Proposition 7.12, Corollary 7.14).
Finally we show that the failure of monotonicity has large cardinal strength, so the large cardinal assumption in Theorem 1.2 cannot be eliminated.
Theorem 1.5 (Theorem 8.4).
@@
Theorem 1.6 (Theorem 8.16).
Suppose there are cardinals $\kappa,\lambda$ such that $\lambda\leq\kappa$ but $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\lambda)$. Then there is an inner model with a proper class of strong cardinals.
@@ 2. Preliminaries
Throughout this paper, a filter means a proper filter over a non-empty set. A filter $\mathcal{F}$ over a set $S$ is uniform if $\left|{X}\right|=\left|{S}\right|$ for every $X\in\mathcal{F}$. For an ultrafilter $\mathcal{U}$ over a set $S$, a subfamily $\mathcal{G}\subseteq\mathcal{U}$ is a base for $\mathcal{U}$ if for every $Y\in\mathcal{U}$ there is $X\in\mathcal{G}$ with $X\subseteq Y$. Let $\chi(\mathcal{U})=\min\{\left|{\mathcal{G}}\right|\mid\mathcal{G}\subseteq\mathcal{U}$ is a base}\}. $\chi(\mathcal{U})$ is called the character of $\mathcal{U}$. Let $\mathfrak{u}(\kappa)$ be the cardinal $\min\{\chi(\mathcal{U})\mid\mathcal{U}$ is a uniform ultrafilter over $\kappa\}$, it is called the ultrafilter number at $\kappa$. It is clear that $\mathfrak{u}(\kappa)\leq 2^{\kappa}$.
Theorem 2.1 (Brendle and Shelah, Proposition 1.4 in [5]).
@@
    There is $f:S\to\lambda$ such that for every $X\subseteq\lambda$ with size $\lambda$, we have $f^{-1}(X)\notin\mathcal{F}$.
@@
Theorem 2.4 (Prikry and Silver [20]).
Let $\lambda<\kappa$ be regular cardinals. If $\kappa$ has a $\lambda$-indecomposable uniform ultrafilter, then every stationary subset of $\{\alpha<\kappa\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$ reflects, that is, for every stationary $S\subseteq\{\alpha<\kappa\mid\mathord{\mathrm{cf}}(\alpha)=\lambda\}$, there is $\beta<\kappa$ such that $\mathord{\mathrm{cf}}(\beta)>\omega$ and $S\cap\beta$ is stationary in $\beta$.
It is known that such a stationary reflection property has large cardinal strength, at least a Mahlo cardinal. See also Inamdar and Rinot [14], and Lambie-Hanson and Rinot [18] for recent developments in this direction.
@@
We note that this observation is immediate by Kunen and Prikry’s result ([17]) that if $\kappa$ is regular, then every $\kappa^{+}$-descendingly incomplete ultrafilter is $\kappa$-descendingly incomplete.
@@
Theorem 2.6 (Ben-David and Magidor [2]).
@@
The following lemma is folklore, and we give a proof of it for the completeness.
@@
Let $\mathcal{U}$ be an ultrafilter over a set $S$. For a set $T$ and $f:S\to T$, let $f_{*}(\mathcal{U})=\{X\subseteq T\mid f^{-1}(X)\in\mathcal{U}\}$. $f_{*}(\mathcal{U})$ is an ultrafilter (possibly principal) over $T$. The next lemma is folklore and can be verified easily.
@@
Let $\mathcal{U}$ be an ultrafilter over a set $S$, and $\lambda$ a cardinal.
  1. (1)
    Let $f:S\to T$. Then $\chi(f_{*}(\mathcal{U}))\leq\chi(\mathcal{U})$, and if $\mathcal{U}$ is $\lambda$-indecomposable, then so is $f_{*}(\mathcal{U})$.
  2. (2)
    $\mathcal{U}$ is $\lambda$-decomposable if and only if there is a function $g:S\to\lambda$ such that $g_{*}(\mathcal{U})$ is a uniform ultrafilter over $\lambda$.
For a regular cardinal $\kappa$ and a non-empty set $X$, let $\mathrm{Add}(\kappa,X)$ be the poset of all partial functions $p:\kappa\times X\to 2$ with size $\kappa$. The order is given by the reverse inclusion. The poset $\mathrm{Add}(\kappa,X)$ is $\kappa$-closed, satisfies the $(2^{<\kappa})^{+}$-c.c., and adds $X$-many new subsets of $\kappa$. For $Y\subseteq X$, we will identify $\mathrm{Add}(\kappa,Y)$ as a complete suborder of $\mathrm{Add}(\kappa,X)$.
@@ 3. Raghavan and Shelah’s theorem
Let $\mathcal{N}=\{\dot{A}\mid\dot{A}$ is a $\gamma$-nice name for some $\gamma<\mu\}$. Since $\mathbb{P}$ has the $\mathord{\mathrm{cf}}(\mu)$-$c.$, $\left|{\{p_{i}\mid i<\gamma\}}\right|<\mu$ for $\gamma<\mu$, and $\nu^{\kappa}<\mu$ for every $\nu<\mu$, we have $\left|{\mathcal{N}}\right|\leq\mu$.
@@
We know that the cardinality of $\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega_{1},\mu)$ is just $\mu$, and has the $\omega_{2}$-c.c. By standard arguments, $\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega_{1},\mu)$ forces that $\kappa=2^{\omega}$ and $2^{\kappa}>2^{\omega_{1}}=\mu$. By Theorem 3.1, we have $\mathfrak{u}(\kappa)\leq\mu<2^{\kappa}$ in the generic extension. ∎
The same argument together with Theorem 2.6 proves:
@@
In the Prikry forcing extension $V^{\mathbb{P}_{\mathcal{U}}}$, it is known that $\mathcal{U}$ generates a $\nu$-indecomposable uniform filter for every regular uncountable $\nu<\kappa$ (Prikry, e.g., see Exercise 18.6 in Kanamori [16]). By Lemma 2.3, every ultrafilter extending $\mathcal{U}$ is $\nu$-indecomposable for regular uncountable $\nu<\kappa$. Hence, in $V^{\mathbb{P}_{\mathcal{U}}}$, $\kappa$ is a singular cardinal with countable cofinality and has a $\mathord{\mathrm{cf}}(\mu)$-indecomposable uniform ultrafilter. Also note that $\mu$ remains strong limit in $V^{\mathbb{P}_{\mathcal{U}}}$. By Theorem 3.1, forcing with $\mathrm{Add}(\omega,\mu)$ over $V^{\mathbb{P}_{\mathcal{U}}}$ forces that $\kappa$ is singular with countable cofinality and $\mathfrak{u}(\kappa)\leq\mu$. Moreover $\mu<2^{\kappa}$ since $\mathord{\mathrm{cf}}(\mu)<\kappa$. ∎